Werner Bley

Title: Equivariant Tamagawa number conjecture for abelian extensions of a quadratic imaginary field.

Abstract: Let k denote a quadratic imaginary field and let L/k be an abelian extension. We will sketch the proof of the p-part of the equivariant Tamagawa number conjecture for the Tate motive h^{0}(Spec(L)) for primes p which split in k/Q.